# Reviewing Intelligent Cinematography: AI research for camera-based video production

### v1

@@ 3.3.2 Object Tracking
[124] presents a lengthy review of ODT approaches including evolutionary algorithms, conventional neural networks and feed-forward neural networks (FNNs). It is shown that there are suitable approaches for almost any combination of known visual constraints. For example, FNNs are most successful for video streams with irregular illumination change and noise, whilst darker hues and low frame rate are best handled by DL-based CNN classifiers.
Within IC we find two distinct paradigms, (1) singular target ODT and (2) multiple target ODT. (1) is much easier to solve and solutions tend to use YOLO detection. For example, Wu et al [270] looks at two sports in particular, swimming and table tennis for over-the-top and in-stadium large-screen broadcasting. With the table tennis case study, problems relate to resolving high-speed motion blur for small moving objects to capture the game live in 3-D. While with swimming, problems relate to light-reflection obscuring the cameras vision of swimmers and water occlusion. For table tennis, a YOLOv4 model is used to estimate ball-bounce from a (singular view) video stream and uses this to define a 3-D trajectory. For swimming the solution is a reduced ResNet-50 neural network model acting as the base-network, alongside a modified SiamRPN++ model for swimmer-tracking. To support recognition the camera’s view is masked to suppress background noise. Wu et al [270] concludes on difficulties with the communication of heterogeneous interfaces between collaborators as well as providing the feasibility of this model in actual practice.
As concluded by Burić et al [32], YOLO methods are the fastest to test and fine-tune [^14] Though, we note only YOLOv2 was tested in this review thus is an attractive case for practical development. They are proven to be sufficient particularly for occlusions. The less successful alternative, Mask R-CNN [93], uses two networks, one for detecting regions of interest (i.e. a Regional Proposal Network (RPN)) and a deep CNN for determining how likely a group of pixels within a region of interest is associated with an object.
@@ 3.3.3 Challenges and Future Work
Relative to other forms of production, live broadcasting has a long standing relationship with AI research. We see this through considerable efforts to elevate the efficacy of pose and object detectors for human targets. We have doubts on the current practicality of these models considering their lack of accuracy relative to human ability, though this is well understood by research and often the main objective. Therefore, we do not expect future work will deviate from investigating this paradigm. Considering the number of processes involved, there are several other feasible avenues for research. For example, we could extend the supporting infrastructure by investigating different masking strategies for suppressing the general set of noisy surroundings. We could also investigate image detection and correction methods of visual artifacts, such as the reflection of light from a swimming pool [270].
Furthermore, a significant sum of footage is readily accessible online. Granted, there are biases in the popularity, gender and racial diversity of sporting events, which should be avoided where possible. Nonetheless, this indicates a need to source more diverse data-sets for general use.
Otherwise, we are interested to see how live IC could play a larger part in driving content capture. For example, we are interested to see more research in optimising shot composition for anticipating events. It would also be exciting to see stylized shot compositions that look at capturing in-game events with a variation of techniques. For instance, using cinematic style transfer [265] to replicate praiseworthy shots. Perhaps one day we may be able to watch a sports game in a Tarantio-esque style.
@@ 3.4 Aerial Production
Similar to many other applications of IC, the problem landscape for UAV-based cinematography is generally undefined [156]. This is further pronounced when we consider the limitations that come with commercial drones, which are not easy to outfit for cinematographic purpose. Mademlis et al [156] produces an interesting review of concurrent challenges, outlining that due to badly defined problem space, lack of technological accessibility as well as roughly drawn legal, ethical and safety constraints, it can be challenging to find a solution which fits the general needs of cinematographers. Despite this, work on autonomous UAVs continues to grow and more recently we find large variety of seemingly successful approaches to UAV control for IC.
In this section we make the key delineation between single drone and multi-drone systems. This is important to consider for cinematographers with tight budgets but also as the technology between the two can greatly differ.
@@ 3.4.1 Automated Single UAV control
Aside from budget constraints, the choice to use a single UAV may come down to the size and accessibility of the camera’s viewing search space. For example a small viewing region around an actor to capture close to mid range shots may only require a single versatile drone. With limited airspace around a subject more drones demands additional constraints to avoid crashes further diminishing the possibility of real-time solutions. Thus there is an area of research on single UAV control referred to as autonomous “follow me” quadrotors [118]. There are a number of vision and sensor-oriented solutions within this domain [173, 246, 145, 60] as well as drone manufacturers that supply this feature [118], like the 3DR Solo and DJI Phantom.
@@
The second challenge to consider is the method of detecting actors. This is useful for trajectory planning but also gimbal control. Bonatti et al [26] considers using a single shot detector [151], such as YOLO9000 [206] and Fast R-CNN [210], for a problem demanding smooth gimbal control and determines that Fast R-CNN is optimal given the speed-accuracy trade off for defining a trajectory. They additionally use MobileNet to perform low-memory feature extraction with fast inference speed. Finally, for indefinite actor tracking they use a KCF [99], taking advantage of the fact that some learning algorithms perform better in the Fourier domain. The KCF tracker relies on kernel ridge regression which is a kernalized version of the linear correlation filter, forming the basis for the fastest trackers available [24, 23].
Finally the third and most evident challenge is how trajectory planning is handled with regard to certain cinematographic objectives. A tempting option is to adapt cinematographic concepts to mathematical expressions that are subsequently optimized for control [195, 118]. Alternatively, Ashtari et al [9] proposes a method that models the dynamic motion of a human camera operator in real time. The authors follow works on modelling the vertical and lateral displacement of a walking patterns [39, 302] combining approaches into a single routine that additionally considers the rotation of the drone and damping effects to simulate different types of camera equipment. Another approach is to let a reinforcement learning agent control the camera motion. Gschwindt et al [88] builds on CHOMP [204] to parameterize smooth trajectory planning while using a deep Q network (DQN) to lead target shot selection. Two methods are proposed for training the DQN. The first method is a human-crafted reward function; like adapting cinematographic shots to optimization functions, this reward function accounts for the actors presences in a shot, the duration and the shot angle. The second method is human-led observation which rewards the DQN relative to the cinematographers subjective opinion.
Ultimately, the design of autonomous single-camera UAV systems is ambiguous. For cinematographers, this means there are a number of trade offs to consider, such as the complexity and feasibility of using more flexibly systems (e.g. drones with gimbals or ”follow me” modes). We agree with the premise set by Mademlis et al [156], that work still needs to be done unifying the objectives of autonomous drone systems. However we also believe that a universal paradigm could inhibit varying uses, for example mimicking human walk patterns compared and optimising shot composition present different objectives and outcomes.
@@ 3.4.3 Challenges and Future Work
Respecting the goal of a general automated UAV solution for IC, the prior subsections lead us to encourage prioritising the discovery of a holistic set solutions. The breadth of use-cases are wide and often unique, however much like the DA problem (discussed in Section3.1.2), solutions heavily rely on the development of related works. For example, solutions to the UAV localization paradigm canonically involve extensions of SLAM and VO research. With recent developments in camera calibration, consequent to the explosion of NeRF research, we are thus likely to see extended use of visual methods for localization and path generation in the future. Considering this avoids the need for GPS-related hardware, it may result in longer flying time and cheaper budgets which are particularly useful for multi-UAV paradigm.
@@
Moreover, with new developments in dynamic and human NeRFs, discussed in Section 3.2.3 and 3.2.4, we see additional potential in training and/or testing automated drones on unforeseen circumstances using real-world digital twins. This could be used to accomplish safety tests or to train UAVs on scene-specific scenarios. For example, if we wanted a UAV to learn to react to actors disobeying drone safety protocol, we could train a NeRF environment using MonoHuman and interpolate sequences of human poses which defy protocol.
@@
On the other hand, the steady developments of NeRF could render UAV-based solutions redundant for directly acquiring the desired content. Rather, it is likely that UAVs become an intermediary step for NeRF-like content acquisition, as it separates the problem into two steps: (1) UAV image acquisition with the objective of optimising a NeRF with 6DoF, and (2) content acquisition within NeRF with 6DoF flexibility. This simplifies the current paradigm, which accomplishes localization and path generation constrained by subjective visual objectives.
@@
Alternatively, while this may apply to the general UAV use-case there are situations that still require attention from IC researchers. For example, the follow-me paradigm implies the use of UAVs for direct content acquisition and relies more on the ability of a drone to move with a visual anchor, like a human target, under cinematographic objectives, like a desired walk pattern or jitter. Considering this is a simpler problem with arguably easier to define objectives, the intervention of NeRF may be unnecessary.
4 General Remarks
This literature review begins with the characterisation of IC as the field of research dedicated to exploring AI solutions for cinematographic video production. We subsequently reviewed four production mediums that we believe are most relevance to current researchers and interested cinematographers, including a subsection on general production tools. This leads us to present our final remarks.
There are certainly several favourable fields of research that have potential to shape the industry going forward:
@@ 4 General Remarks
    Computational Language Structures, discussed in Section 3.1.1
@@
    LED Volumes, discussed in Section 3.2.1
@@
    NeRFs, discussed in Section 3.2.3 and 3.2.4
With regards to computational language structures, we do not believe that the current state of research is telling of its potential impact on IC. Instead, we believe with the current state of research on LLMs and it’s use in the public domain [^17] E.g. generating trailers from text/scripts [211, 54], that natural language decision making for IC will be facilitated. Though we have yet to see this is investigated in research, we believe there is cause to pursue its application in fields like automated directive assistance, MP for planning and automated UAV control.
With regards to LED volumes and NeRFs we believe the impact on industry has already been felt. For the former, the adoption of this new technology has offered radical change to canonical chroma-keying practices and offers a new image-based paradigm for computer vision researchers. Regarding the latter, research has demonstrated a flexible pool of new solutions to problems which were previously solved through with sculpting/3-D modelling tools and/or hardware dependant photogrammetric methods. With the current push for compact, fast and accurate representations, we would not be surprised if production ready NeRFs became available soon.
Conclusively, we hope that our work inspires more reviews on future developments of IC research. The fields of research and the relevant video production industries we have discussed are highly susceptible to change, as exemplified by the influence of the LED volume. Thus we should acknowledge the need for concurrent reviews. Finally, to further bridge the gap between IC researchers and industry professionals we encourage future reviews that targets industry professionals directly.
4.1 Social Responsibility
The impact of evolving technologies on the creative industries is multifaceted. Regarding IC, there are two matters that underpin widespread concerns with freely advancing AI: (1) replacing actors with human-like AI models [17, 51, 52] and (2) using generative AI and/or 3-D models to produce pornographic content [245, 178]. The matters mainly concern fields such as automated 3-D capture and generative modelling, where the ability to capture a persons physical essence is currently possible to a high degree for images and likely soon-to-be for videos. As we discuss these topics in Section 3.2.4 and 2.2, respectively, we take the opportunity now to provide awareness on the issues and discuss steps which can be taken to avoid certain ethical dilemmas.
Replacing real actors can be achieved by either modelling a real human or generating a virtual human [17]. For actors this is problematic as current stars fear abuse of existing content containing detailed depictions of themselves. This may also lead to marginalizing real people for virtual celebrity look-a-likes.

### v3

@@ 3.3.2 Object Tracking
[133] presents a comprehensive review of ODT approaches including evolutionary algorithms, conventional neural networks and feed-forward neural networks (FNNs). It is shown that there are suitable approaches for almost any combination of known visual constraints. For example, FNNs are most successful for video streams with irregular illumination change and noise, whilst darker hues and low frame rate are best handled by DL-based CNN classifiers.
Within IC we find two distinct paradigms, (1) singular target ODT and (2) multiple target ODT. (1) is much easier to solve and solutions tend to use YOLO detection. For example, Wu et al [283] looks at two sports in particular, swimming and table tennis for over-the-top and in-stadium large-screen broadcasting. With the table tennis case study, problems relate to resolving high-speed motion blur for small moving objects to capture the game live in 3-D. While with swimming, problems relate to light-reflection obscuring the cameras vision of swimmers and water occlusion. For table tennis, a YOLOv4 model is used to estimate ball-bounce from a (singular view) video stream and uses this to define a 3-D trajectory. For swimming the solution is a reduced ResNet-50 neural network model acting as the base-network, alongside a modified SiamRPN++ model for swimmer-tracking. To support recognition the camera’s view is masked to suppress background noise. Wu et al [283] concludes on difficulties with the communication of heterogeneous interfaces between collaborators as well as providing the feasibility of this model in practice.
As concluded by Burić et al [33], YOLO methods are the fastest to test and fine-tune [^12] We note only YOLOv2 was tested in this review and thus are an attractive candidate for practical development. They have also been proven to work well in the presence of occlusions. The less successful alternative, Mask R-CNN [99], uses two networks, one for detecting regions of interest (i.e. a Regional Proposal Network (RPN)) and a deep CNN for determining how likely a group of pixels within a region of interest is associated with an object.
@@ 3.3.3 Challenges and Future Work
Relative to other forms of production, live broadcasting has a long standing relationship with AI research. For example, the BBC have been involved with supporting AI research in live production settings, by collecting and freely providing the Old School dataset [126] to support televised production tasks such as enhanced shot framing, [125]. The relationship with industry can be seen through the considerable efforts to elevate the efficacy of pose and object detectors for human targets. However, We have doubts on the current practicality of these models considering their lack of accuracy relative to human ability, although this is well understood by the research community. Considering the number of processes involved, there are several other feasible avenues for research. For example, we could extend the supporting infrastructure by investigating different masking strategies for suppressing the general set of noisy surroundings. We could also investigate image detection and correction methods of visual artifacts, such as the reflection of light from a swimming pool [283].
Furthermore, a significant quantity of footage is readily accessible online. However researchers must be cognizant of biases in the popularity, gender and racial diversity of sporting events. This indicates a need to source more diverse data-sets for general use.
IC could play a larger part in driving content capture. For example, more research in optimizing shot composition for anticipating events would be of benefit. There may also be interest in stylized shot compositions that present in-game events, for instance, using cinematic style transfer [277] to emphasize praiseworthy shots. Perhaps one day we may be able to watch a sports game in a Tarantio-esque style.
@@ 3.4 Aerial Production
Similar to many other applications of IC, the problem landscape for UAV-based cinematography is generally poorly defined [166]. Mademlis et al [166] produces an interesting review of concurrent challenges, outlining that due to badly defined problem space, lack of technological accessibility as well as roughly drawn legal, ethical and safety constraints, it can be challenging to find a solution which fits the general needs of cinematographers.
In this section we describe both single drone and multi-drone systems. This is important to consider for cinematographers with tight budgets but also as their technology, capability and complexity differs significantly.
@@ 3.4.1 Automated Single UAV control
Aside from budget constraints, the choice to use a single UAV may come down to the size and accessibility of the camera’s viewing space. For example a small viewing region around an actor to capture close to mid range shots may only require a single versatile drone. With limited airspace around a subject more drones demands additional constraints to avoid crashes further diminishing the possibility of real-time solutions. Thus there is an area of research on single UAV control referred to as autonomous “follow me” quadrotors [127]. There are a number of vision and sensor-oriented solutions within this domain [183, 258, 155, 62] as well as drone manufacturers that supply this feature [127], like the 3DR Solo and DJI Phantom.
@@
The second challenge is the method of detecting actors. This is useful for trajectory planning but also for gimbal control. Bonatti et al [26] considers using a single shot detector [161], such as YOLO9000 [217] and Fast R-CNN [221], for a problem demanding smooth gimbal control and determines that Fast R-CNN is optimal given the speed-accuracy trade off for defining a trajectory. They additionally use MobileNet to perform low-memory feature extraction with fast inference speed. Finally, for indefinite actor tracking they use a KCF [105], taking advantage of the fact that some learning algorithms perform better in the Fourier domain. The KCF tracker relies on kernel ridge regression which is a kernalized version of the linear correlation filter, forming the basis for the fastest trackers available [24, 23].
Finally the third and most evident challenge is how trajectory planning is handled with regard to certain cinematographic objectives. A tempting option is to adapt cinematographic concepts to mathematical expressions that are subsequently optimized for control [206, 127]. Alternatively, Ashtari et al [8] proposes a method that models the dynamic motion of a human camera operator in real time. The authors follow works on modeling the vertical and lateral displacement of a walking patterns [41, 315] combining approaches into a single routine that additionally considers the rotation of the drone and damping effects to simulate different types of camera equipment. Another approach is to let a reinforcement learning agent control the camera motion. Gschwindt et al [94] builds on CHOMP [215] to parameterize smooth trajectory planning while using a deep Q network (DQN) to lead target shot selection. Two methods are proposed for training the DQN. The first method is a human-crafted reward function; like adapting cinematographic shots to optimization functions, this reward function accounts for the actors presences in a shot, the duration and the shot angle. The second method is human-led observation which rewards the DQN relative to the cinematographers subjective opinion.
Ultimately, the design of autonomous single-camera UAV systems is ambiguous. For cinematographers, this means there are a number of trade offs to consider, such as the complexity and feasibility of using more flexibly systems (e.g., drones with gimbals or ”follow me” modes). We agree with the premise set by Mademlis et al [166], that work still needs to be done unifying the objectives of autonomous drone systems. However we also believe that a universal paradigm could inhibit varying uses, for example mimicking human walk patterns compared and optimizing shot composition present different objectives and outcomes.
@@ 3.4.3 Challenges and Future Work
As with the live production setting, aerial cinematography has been strongly influence technological developments. For instance, AI-controlled drone cinematography is used to capture live events by organizations such as Quest Films [^15] More accessible here: https://quest-films.com [205]. Even commercially available drones are now equipped with AI capture and control features, such the follow me mode in Altair, DJI and Hubsan products.
@@
Following [166], we emphasize the need for discovering a holistic set solutions, building on the wider research literature. For example, solutions to the UAV localization paradigm canonically involve extensions of SLAM and VO research. With recent developments in camera calibration, consequent to the explosion of NeRF research, we are thus likely to see extended use of visual methods for localization and path generation in the future. Considering this avoids the need for GPS-related hardware, it may result in longer flying time and cheaper budgets which are particularly useful for the multi-UAV paradigm.
@@
Moreover, with new developments in dynamic and human NeRFs, discussed in Section 3.2.3 and 3.2.4, we see significant potential in training and/or testing automated drones on unforeseen real-world circumstances using digital twins. Boyle et al [27] investigates the use of photogrammetry for training camera operators on single-opportunity shots in live event broadcasting. While the breadth and complexity of test scenarios are limited, we believe NeRFs have potential to fill the gaps, such as introducing more complex human-led interactions. This could also be used to accomplish safety tests so that UAVs, for example, may learn to react to actors disobeying drone safety protocol.
@@
On the other hand, the steady developments in the field of of NeRF technologies could render UAV-based solutions redundant for directly acquiring actor-centric content. Rather, it is likely that UAVs become an intermediary step for NeRF-like content acquisition, as it separates the problem into two steps: (1) UAV image acquisition with the objective of optimizing a NeRF with 6DoF, and (2) content acquisition within NeRF with 6DoF flexibility. This simplifies the current paradigm, which accomplishes localization and path generation constrained by subjective visual objectives.
@@
Alternatively, while this may apply to the general UAV use-case, there are situations that still require attention from IC researchers. For example, the follow-me paradigm implies the use of UAVs for direct content acquisition and relies more on the ability of a drone to move with a visual anchor, like a human target, under cinematographic objectives, like a desired walk pattern or jitter. Considering this is a simpler problem with arguably easier to define objectives, the intervention of NeRF may be unnecessary for specific applications.
3.5 Social, Legal and Ethical Challenges
The impact of evolving technologies, in particular AI, on the creative industries is multifaceted and fast changing. In the context of IC, there are several issues that underpin widespread concerns regarding the negative impact of recent AI breakthroughs, such as:
@@ 3.5 Social, Legal and Ethical Challenges
    Replacing actors with human-like AI models [16, 53, 54]
@@
    Using generative AI and/or 3-D models to produce pornographic content [257, 188]
@@
    Using existing art to train generative AI
In the following subsections we expand on these concerns as well as highlight grounded counter arguments where applicable. There are a large number of perspectives to consider, which can be influenced by things like the degree of personal involvement (e.g., researchers involved in generative AI, producers struggling to meet budgets or artists competing against AI for work, etc.), trust in certain sources of information (e.g., news organizations or popular creatives voicing their opinions) or even the spread of mis-information. Hence, we have tried to remain impartial on these complex topics and encourage readers to form their own opinion.
3.5.1 Replacing Actors with AI Avatars
Replacing real actors can be achieved by either modeling a real human or generating a virtual human [16]. For actors this can be problematic as they fear abuse or exploitation of existing content containing detailed depictions of themselves. This may also lead to marginalizing real people for virtual celebrity look-a-likes. These points have been emphasized by SAG-AFTRA, a union that represent 160,000 media professionals and was a primary reason for the SAG-AFTRA and Writers Guilds of America strikes that began in 2023. The proposed regulations for AI provided by [229] primarily deal with consent and compensation by essentially demanding that digital replicas be treated the same as the real media professionals, i.e. “Residuals paid for use that would normally generate residuals”. Digital alteration is also discussed where the need for reasoning (for making cosmetic modification) and transparency (details of the alterations should be conveyed to the union) is emphasized. Moving forward, we believe researchers should consider the ethical implications of producing work that relates to the generation or modification of digital twins or virtual avatars. When work may potentially impact workers, we encourage demonstrating awareness by being transparent about the social impact of future use cases. Importantly, considering the potential future legal ramifications for researchers, more ethical practice could manifest by monitoring and publicly disclosing the use and development of AI datasets and tools as proposed by European Commission, [78].
3.5.2 AI Generated Pornography